21 June 2012
The Sun reappeared!
Hurrah, after a week of overcast we saw the Sun today!
Both wind speeds and outdoor temperatures have been dropping, so we are in a better shape to prepare for next week’s flight and there was much activity on the surface: Lou started preparing the apron, Sverrir removed snow dunes in camp, Lars and Nanna went out to measure a couple of GPS positions of the NEEM strain net, and the science trench gang moved another pile of gear to surface. If those conditions continue, we may soon try opening the ramps to the trenches and buildings again.
What we have done today:
1. No drilling today
2. NEGIS: wind around 10 kts, the Sun is seen, and there is good progress
3. Redistributing large amounts of snow in camp
4. Started beam grooming apron
5. Measuring two GPS stations at 7.5 km and 25 km
6. Packing of science trench electronics
7. Moving science trench stuff to pallets
8. Worked on synchronization of NEEM and NEGIS ice cores
Ad. 1: The deep drilling cannot proceed any further without more extension rods for the rock drill. It was somewhat unexpected that the drilling has not yet reached ‘solid’ bedrock at this point. There is general agreement that the drilling should continue if possible. Therefore, new rods have been ordered and the plan is to continue the deep drilling in 2-3 weeks time.
Ad. 6: The science trench is slowly being emptied. It now contains ice core boxes, the cooks freezer, a seismic station and some bulky items that need opening of the drill trench ramp to get out.
Weather: Mostly overcast and snow drift most of the day, but sunny spells appeared. Wind 10-30 knots from SW, rapidly decreasing at night. Temp. -3°C to -10°C, decreasing at night.
FL, Anders Svensson